OpenStudy (anonymous):

A source of laser light sends rays AB and AC toward two opposite walls of a hall. The light rays strike the walls at points B and C, as shown below.

OpenStudy (anonymous):

please do NOT just tell me the answer I want an explanation

OpenStudy (anonymous):

What is the distance between the walls?

OpenStudy (anonymous):

Let the remaining vertices of the triangles with vertex B and vertex C be given by D and E respectively: Hence in triangle ABD, we have : \[ \huge \cos \theta = \frac{DA}{AB} \rightarrow \cos 60^0 = \frac{DA}{40}\] \[ \huge \rightarrow \frac{1}{2} = \frac{DA}{40} \rightarrow DA = 20 m\] Next in triangle ACE, we have \[\huge \tan \theta = \frac{CE}{AE} \rightarrow \tan 45^0 = \frac{80}{AE}\] \[\huge \rightarrow 1 = \frac{80}{AE} \rightarrow AE = 80 m\] Now distance between two walls = DA + AE = 20 + 80 = 100 m @Sashamichelle

OpenStudy (anonymous):

thanks so much!